# TACLR: A Scalable and Efficient Retrieval-based Method for Industrial Product Attribute Value Identification

### v1

@@ Abstract
Product Attribute Value Identification (PAVI) involves identifying attribute values from product profiles, a key task for improving product search, recommendations, and business analytics on e-commerce platforms. However, existing PAVI methods face critical challenges, such as inferring implicit values, handling out-of-distribution (OOD) values, and producing normalized outputs. To address these limitations, we introduce Taxonomy-Aware Contrastive Learning Retrieval (TACLR), the first retrieval-based method for PAVI. TACLR formulates PAVI as an information retrieval task by encoding product profiles and candidate values into embeddings and retrieving values based on their similarity to the item embedding. It leverages contrastive training with taxonomy-aware hard negative sampling and employs adaptive inference with dynamic thresholds. TACLR offers three key advantages: (1) it effectively handles implicit and OOD values while producing normalized outputs; (2) it scales to thousands of categories, tens of thousands of attributes, and millions of values; and (3) it supports efficient inference for high-load industrial scenarios. Extensive experiments on proprietary and public datasets validate the effectiveness and efficiency of TACLR. Moreover, it has been successfully deployed in a real-world e-commerce platform, processing millions of product listings daily while supporting dynamic, large-scale attribute taxonomies.
@@ 1 Introduction
Figure 1: Illustration of the PAVE and PAVI tasks. While an additional normalization step can adapt PAVE methods for PAVI, these methods remain unable to identify implicit values, such as 6GB.
Product attribute values are critical components that underpin the functionality of e-commerce platforms. They provide essential structural information, aiding customers in making informed purchasing decisions and enabling various services, such as product listing Chen et al. (2024), recommendation Truong et al. (2022); Sun et al. (2020), retrieval Huang et al. (2014), and question answering Kulkarni et al. (2019); Gao et al. (2019).
However, seller-provided attribute values are often incomplete or inaccurate, undermining the effectiveness of these applications. As a result, the automatic identification of product attribute values has become a critical research challenge.
Researchers have explored Product Attribute Value Extraction (PAVE), which extracts spans from product profiles using Named Entity Recognition (NER) Zheng et al. (2018) or Question Answering (QA) Wang et al. (2020) models. The upper part of Figure 1 illustrates an example of NER-based PAVE. While this paradigm effectively extracts value spans, the outputs are raw subsequences. To produce standardized values, a normalization step Putthividhya and Hu (2011); Zhang et al. (2021) is required to map these spans to predefined formats, as shown in the lower part of Figure 1. However, implicit values, such as 6GB, cannot be directly extracted and must instead be inferred from context or prior knowledge.
Presenting attribute values in a standardized format is crucial for facilitating data aggregation in business analytics and enhancing the user experience by providing clear and consistent information. Therefore, in this work, we focus on the more challenging task of Product Attribute Value Identification (PAVI) Shinzato et al. (2023), which aims to associate predefined attribute values from attribute taxonomy (illustrated in Figure 2) with products. The input to PAVI includes the product category and profile, where the profile comprises textual data, such as the title and description, and may optionally include visual information, such as images or videos. The output is a dictionary with predefined attributes as keys and the inferred attribute values as corresponding entries. In addition, PAVI requires determining when attribute values are missing. For instance, as shown in Figure 1, the value for Version is unavailable and is therefore assigned an empty result or null value.
@@ 2.1 Product Attribute Value Extraction
PAVE as Named Entity Recognition. PAVE can be formulated as NER by identifying subsequences in product texts as entity spans and linking them to attributes as entity types. Early methods, such as OpenTag Zheng et al. (2018), trained individual models for each category-attribute pair. Subsequent efforts generalized this approach to support multiple attributes or categories. For instance, SUOpenTag Xu et al. (2019) incorporated attribute embeddings into an attention layer to handle multiple attributes, while AdaTag Yan et al. (2021) used attribute embeddings to parameterize the decoder. TXtract Karamanolakis et al. (2020) introduced a category encoder and a category attention mechanism to tackle various categories effectively. Additionally, M-JAVE Zhu et al. (2020) jointly modeled attribute prediction and value extraction tasks while also incorporating visual information. More recently, Chen et al. (2023) scaled BERT-NER by expanding the number of entity types to support a broader range of attributes.
@@ 3.2 Retrieval-Based PAVI
To effectively encode both the item and candidate values, we preprocess them as textual inputs and use a shared text encoder. For each item, we concatenate its title $t$ and description $d$ as the input text. Each candidate value $v$ under an attribute $a$ within category $c$ is represented as a text prompt, such as “A [phone ($c$)] with [brand ($a$)] being [Apple ($v$)]”.  [^1] This framework can be extended to multimodal scenarios by replacing the text encoder with a multimodal encoder to incorporate features like images. We explore the impact of various value prompt templates in §5.3.
@@ 3.4 Adaptive Inference
To address this, we introduce an adaptive inference method that uses dynamic thresholds to make cutoff decisions. As discussed in §3.3, we add an explicit null value $v_{0}^{a}$ for each category-attribute pair, with its embedding learned during training. In the inference phase, we compute the similarity $s(i,v_{0}^{a})$ between the item and the null value, using it as a dynamic threshold $T^{\prime}_{a}$ to filter out values with scores lower than that of the null value:
@@
Since most category-attribute pairs have exclusive values, meaning that each product can have at most one value for a given attribute, we focus on the top 1 predicted value in this work. The output can be further simplified as follows:
@@ 4.2 Metrics
Table 4: Performance comparison of different methods on Ecom-PAVI and WDC-PAVE. "F1 Excl." refers to the F1 score calculated without numeric attributes (e.g., width and height), which require unit conversion reasoning.
@@ 4.3 Baselines
We evaluate our retrieval-based method TACLR against classification and generation baselines. [^4] Extraction-based baselines, such as NER or QA models, are widely used for PAVE, but they are excluded from our comparison due to the lack of standard normalization steps, which makes fair evaluation challenging. For implementation details, refer to Appendix A.
@@ 5 Results
BERT-CLS 8.6 93.1
Llama3.1 (zero-shot) 101.3 7.9
Llama3.1 (few-shot) 124.8 6.4
Qwen2.5 (zero-shot) 84.0 9.5
Qwen2.5 (few-shot) 98.4 9.1
TACLR 12.7 63.2
@@ 5.1 Main Results
Table 4 presents the performance comparison between our retrieval-based method TACLR and classification- and generation-based baselines on Ecom-PAVI and WDC-PAVE. On Ecom-PAVI, TACLR achieves the highest F1 score of 86.2, surpassing the fine-tuned Llama3.1, which obtains an F1 score of 84.7. Notably, TACLR excels in recall, achieving 87.1 compared to Llama3.1’s 82.7. On WDC-PAVE, TACLR achieves the highest F1 Excl. score of 80.3, which excludes measurement attributes requiring reasoning ability for unit normalization. This result highlights TACLR’s effectiveness and robustness in addressing general PAVI across diverse datasets.
The classification-based method, BERT-CLS, shows the weakest performance on both datasets. It achieves an F1 score of 50.5 on Ecom-PAVI, but its performance drops drastically on WDC-PAVE, where it only attains an F1 score of 20.5. This underscores the limitations of classification-based approaches in generalization, including their inability to adapt to OOD values.
Among the generation-based methods, few-shot and fine-tuning consistently improve performance over zero-shot settings. For example, on Ecom-PAVI, Llama3.1 achieves F1 scores of 35.7, 38.6, and 84.7 in zero-shot, few-shot, and fine-tuned settings, respectively. Similarly, Qwen2.5 achieves F1 scores of 48.4, 51.4, and 81.7 in the corresponding settings. On WDC-PAVE, however, fine-tuned LLMs exhibit weaker generalization due to the scarcity of training data. Few-shot learning proves more robust in this dataset, achieving an F1 score of 75.0 for Llama3.1 and 72.2 for Qwen2.5.
@@ 5.2 Inference Efficiency
TACLR achieves a strong balance between performance and efficiency, with an inference time of 12.7 ms and a throughput of 63.2 samples per second. In contrast, generation-based methods, such as Llama3.1 (few-shot) and Qwen2.5 (few-shot), exhibit significantly longer inference times (124.8 ms and 98.4 ms, respectively) and lower throughputs (6.4 and 9.1 samples per second), highlighting the computational overhead of LLMs in high-load scenarios. While BERT-CLS delivers the fastest inference time and highest throughput, its inability to handle OOD values and limited capacity restrict its effectiveness in practical applications.
@@ 5.3 Analysis
Comparison of Dynamic and Static Thresholds. Figure 5(a) evaluates the dynamic, learnable thresholds (§3.4) against the best static thresholds of 0.6, 0.65, and 0.7, which were chosen based on their performance on the validation set. The dynamic threshold achieves the highest F1 score of 86.2%, surpassing the static thresholds, which yield F1 scores of 75.5%, 80.2%, and 84.5%, respectively. Static thresholds exhibit a clear trend: as the threshold increases, precision rises (from 74.8% to 87.1%) while recall diminishes (from 86.3% to 72.8%). In contrast, the dynamic threshold balances precision (85.4%) and recall (87.1%) effectively, eliminating the need for extensive hyperparameter tuning across category-attribute pairs. This adaptability makes dynamic thresholds a practical choice.
@@
Zero-Shot Generalization Across Data Domains. Figure 5(c) presents the results of zero-shot transfer experiments, evaluating the model’s generalization across unseen categories and values. The in-domain split achieves an F1 score of 88.7%, while performance decreases for cross-category and cross-value splits, which attain F1 scores of 80.2% and 78.2%, respectively. These declines reflect the inherent challenges of adapting to dynamic attribute taxonomies in OOD domains. Nevertheless, the overall F1 score of 85.4% demonstrates robust generalization capabilities.
@@ 6 Conclusion
In this work, we present TACLR, a novel approach for retrieval-based PAVI. By formulating PAVI as an information retrieval task, TACLR inherently supports generalization to OOD values, infers implicit values, and produces normalized outputs. Building on this framework, TACLR employs contrastive training with taxonomy-aware sampling and adaptive inference with dynamic thresholds to enhance retrieval performance and scalability.
Comprehensive experiments on proprietary and public datasets demonstrated TACLR’s superiority over classification- and generation-based baselines. Notably, TACLR achieved an F1 score of 86.2% on the large-scale Ecom-PAVI dataset. Efficiency analysis further highlighted its advantage, achieving significantly faster inference speeds than generation-based methods. TACLR has been successfully deployed on a real-world e-commerce platform, processing millions of product listings daily and seamlessly adapting to dynamic attribute taxonomies, making it a practical solution for large-scale industrial applications.
@@ Appendix B Deployment
In the product listing process, TACLR automatically identifies attribute-value pairs from user-provided titles and descriptions, significantly reducing manual input and errors while improving the quality of structured information.
For product search and recommendation, the enhanced structured data provided by TACLR enables more accurate item retrieval, improving product matching with user queries and personalized recommendations.

### v2

@@ Abstract
Product Attribute Value Identification (PAVI) involves identifying attribute values from product profiles, a key task for improving product search, recommendation, and business analytics on e-commerce platforms. However, existing PAVI methods face critical challenges, such as inferring implicit values, handling out-of-distribution (OOD) values, and producing normalized outputs. To address these limitations, we introduce Taxonomy-Aware Contrastive Learning Retrieval (TACLR), the first retrieval-based method for PAVI. TACLR formulates PAVI as an information retrieval task by encoding product profiles and candidate values into embeddings and retrieving values based on their similarity to the item embedding. It leverages contrastive training with taxonomy-aware hard negative sampling and employs adaptive inference with dynamic thresholds. TACLR offers three key advantages: (1) it effectively handles implicit and OOD values while producing normalized outputs; (2) it scales to thousands of categories, tens of thousands of attributes, and millions of values; and (3) it supports efficient inference for high-load industrial deployment. Extensive experiments on proprietary and public datasets validate the effectiveness and efficiency of TACLR. Moreover, it has been successfully deployed in a real-world e-commerce platform, processing millions of product listings daily while supporting dynamic, large-scale attribute taxonomies.
@@ 1 Introduction
Figure 1: Illustration of the PAVE and PAVI tasks. While an additional normalization step can adapt PAVE methods for PAVI, these methods remain unable to identify implicit values, such as Apple.
Product attribute values are critical components that support the function of e-commerce platforms. They provide essential structural information, aiding customers in making informed purchasing decisions and enabling product listing Chen et al. (2024), recommendation Truong et al. (2022); Sun et al. (2020), retrieval 10.1145/3308560.3316603; Huang et al. (2014), and question answering Kulkarni et al. (2019); Gao et al. (2019).
However, seller-provided attribute values are often incomplete or even inaccurate. This undermines the effectiveness of applications that rely on this information. Consequently, the automatic identification of product attribute values has become a critical challenge. Researchers have explored the task of Product Attribute Value Extraction (PAVE), which extracts spans from product profiles using Named Entity Recognition (NER) Zheng et al. (2018) or Question Answering (QA) Wang et al. (2020) models. The upper part of Figure 1 illustrates an example of NER-based PAVE.
Although these approaches effectively extract value spans, the outputs remain raw subsequences. Presenting attribute values in a standardized format is crucial for facilitating data aggregation in business analytics and enhancing the user experience by providing clear and consistent information. To produce standardized values, a normalization step Putthividhya and Hu (2011); Zhang et al. (2021) is required to map these spans to predefined formats, as shown in the lower part of Figure 1. However, implicit values, such as Apple, cannot be directly extracted and must instead be inferred from context or prior knowledge.
Therefore, in this work, we focus on the task of Product Attribute Value Identification (PAVI) Shinzato et al. (2023), which aims to associate predefined attribute values from attribute taxonomy (illustrated in Figure 2) with products. The input to PAVI includes the product category and profile, where the profile comprises textual data, such as the title and description, and may optionally include visual information, such as images or videos. The output is a dictionary with predefined attributes as keys and the inferred attribute values as corresponding entries. In addition, PAVI requires determining when attribute values are missing. For instance, as shown in Figure 1, the value for Version is unavailable and is therefore assigned an empty result or null value.
@@ 2.1 Product Attribute Value Extraction
PAVE as Named Entity Recognition. PAVE can be formulated as NER by identifying subsequences in product texts as entity spans and associating them with attributes as entity types. Early methods, such as OpenTag Zheng et al. (2018), trained individual models for each category-attribute pair. Subsequent efforts generalized this approach to support multiple attributes or categories. For instance, SUOpenTag Xu et al. (2019) incorporated attribute embeddings into an attention layer to handle multiple attributes, while AdaTag Yan et al. (2021) used attribute embeddings to parameterize the decoder. TXtract Karamanolakis et al. (2020) introduced a category encoder and a category attention mechanism to tackle various categories effectively. Additionally, M-JAVE Zhu et al. (2020) jointly modeled attribute prediction and value extraction tasks while also incorporating visual information. More recently, Chen et al. (2023) scaled BERT-NER by expanding the number of entity types to support a broader range of attributes.
@@ 3.2 Retrieval-Based PAVI
To effectively encode both the item and candidate values, we preprocess them as textual inputs and use a shared text encoder. For each item, we concatenate its title $t$ and description $d$ as the input text. Each candidate value $v$ associated with attribute $a$ under category $c$ is represented as a text prompt, such as “A [phone ($c$)] with [brand ($a$)] being [Apple ($v$)]”.  [^1] This framework can be extended to multimodal scenarios by replacing the text encoder with a multimodal encoder to incorporate features like images. We explore the impact of various value prompt templates in §5.3.
@@ 3.4 Adaptive Inference
To address this, we introduce an adaptive inference method that uses dynamic thresholds to make cutoff decisions. As discussed in §3.3, we add an explicit null value $v_{0}^{a}$ for each category-attribute pair, with its embedding learned during training. In the inference phase, we compute the similarity $s(i,v_{0}^{a})$ between the item and the null value, using it as a dynamic threshold $T^{\prime}_{a}$ to exclude candidate values for attribute $a$ that have lower scores:
@@
Since most category-attribute pairs have exclusive values, meaning that each product can have at most one value for a given attribute, we focus on the top-1 predicted value in this work. The output can be further simplified as follows:
@@ 4.2 Metrics
Table 4: Performance comparison of different methods on Ecom-PAVI and WDC-PAVE. "F1 Excl." refers to the F1 score calculated without measurement attributes (e.g., width and height), which require unit conversion reasoning.
@@ 4.3 Baselines
We evaluate our retrieval-based method TACLR against classification and generation baselines. [^4] Extraction-based baselines, such as NER or QA models, are widely used for PAVE, but they are excluded from our comparison due to the lack of a standard normalization method, which makes fair evaluation challenging. For implementation details, refer to Appendix A.
@@ 5 Results
BERT-CLS 8.6 930
Llama3.1 (zero-shot) 101.3 80
Llama3.1 (few-shot) 124.8 64
Qwen2.5 (zero-shot) 84.0 95
Qwen2.5 (few-shot) 98.4 81
TACLR 12.7 630
@@ 5.1 Main Results
Table 4 presents the performance comparison between our retrieval-based method TACLR and classification- and generation-based baselines on Ecom-PAVI and WDC-PAVE. On Ecom-PAVI, TACLR achieves the highest F1 score of 86.2%, surpassing the fine-tuned Llama3.1, which obtains an F1 score of 84.7%. Notably, TACLR excels in recall, achieving 87.1% compared to Llama3.1’s 82.7%. On WDC-PAVE, TACLR achieves the highest F1 Excl. score of 80.3%, which excludes measurement attributes requiring reasoning ability for unit normalization. This result highlights TACLR’s effectiveness and robustness in addressing general PAVI across diverse datasets.
The classification-based method, BERT-CLS, shows the weakest performance on both datasets. It achieves an F1 score of 50.5% on Ecom-PAVI, but its performance drops drastically on WDC-PAVE, where it only attains an F1 score of 20.5%. This underscores the limitations of classification-based approaches in generalization, including their inability to adapt to OOD values.
Among the generation-based methods, few-shot and fine-tuning consistently improve performance over zero-shot settings. For example, on Ecom-PAVI, Llama3.1 achieves F1 scores of 35.7%, 38.6%, and 84.7% in zero-shot, few-shot, and fine-tuned settings, respectively. Similarly, Qwen2.5 achieves F1 scores of 48.4%, 51.4%, and 81.7% in the corresponding settings. On WDC-PAVE, however, fine-tuned LLMs exhibit weaker generalization due to the scarcity of training data. Few-shot learning proves more robust in this dataset, achieving an F1 score of 75.0% for Llama3.1 and 72.2% for Qwen2.5.
@@ 5.2 Inference Efficiency
TACLR achieves a strong balance between performance and efficiency, with an inference time of 12.7 ms and a throughput of 630 samples per second. In contrast, generation-based methods, such as Llama3.1 (few-shot) and Qwen2.5 (few-shot), exhibit significantly longer inference times (124.8 ms and 98.4 ms, respectively) and lower throughputs (64 and 81 samples per second), highlighting the computational overhead of LLMs in high-load scenarios. While BERT-CLS delivers the fastest inference time and highest throughput, its inability to handle OOD values and limited capacity restrict its effectiveness in practical applications.
@@ 5.3 Analysis
Comparison of Dynamic and Static Thresholds. Figure 5(a) evaluates the dynamic, learnable thresholds (§3.4) against the best static thresholds of 0.6, 0.65, and 0.7, which were chosen based on their performance on the validation set. The dynamic threshold achieves the highest F1 score of 86.2%, surpassing the static thresholds, which yield F1 scores of 75.5%, 80.2%, and 78.2%, respectively. Static thresholds exhibit a clear trend: as the threshold increases, precision rises (from 65.1% to 84.5%) while recall diminishes (from 90.0% to 72.8%). In contrast, the dynamic threshold balances precision (85.4%) and recall (87.1%) effectively, eliminating the need for extensive hyperparameter tuning across category-attribute pairs. This adaptability makes dynamic thresholds a practical choice.
@@
Zero-Shot Generalization Across Data Domains. Figure 5(c) presents the results of zero-shot transfer experiments, evaluating the model’s generalization across unseen categories and values. The in-domain split achieves an F1 score of 88.7%, while performance decreases for cross-category and cross-value splits, which attain F1 scores of 80.2% and 78.2%, respectively. These declines reflect the inherent challenges of adapting to dynamic attribute taxonomies in OOD domains. Nevertheless, the overall F1 score of 86.2% demonstrates the robust generalization capabilities of TACLR.
@@ 6 Conclusion
In this work, we present TACLR, a novel approach for retrieval-based PAVI. By formulating PAVI as an information retrieval problem, TACLR enables the inference of implicit values, generalization to OOD values, and the production of normalized outputs. Building on this framework, TACLR employs contrastive training with taxonomy-aware sampling and adaptive inference with dynamic thresholds to enhance retrieval performance and scalability.
Comprehensive experiments on proprietary and public datasets demonstrated TACLR’s superiority over classification- and generation-based baselines. Notably, TACLR achieved an F1 score of 86.2% on the large-scale Ecom-PAVI dataset. Efficiency analysis further highlighted its advantage, achieving significantly faster inference speeds than generation-based methods. Beyond these experimental results, TACLR has been successfully deployed on a real-world e-commerce platform, processing millions of product listings daily and seamlessly adapting to dynamic attribute taxonomies, making it a practical solution for large-scale industrial applications.
@@ Appendix B Deployment
In the product listing process, TACLR automatically identifies attribute-value pairs from user-provided titles and descriptions, significantly reducing manual effort and errors while improving the quality of structured information.
For product search, improved structured information directly enhances lexical retrieval, leading to more accurate matching with user queries. Meanwhile, it enriches product features, consequently enhancing personalized recommendations.